OpenStudy (anonymous):

Darcy McMillan borrowed $3,600 for 7 months nd had an interest charge of $250. Find the approximate annual interest rate to the nearest tenth of a percent.

OpenStudy (amistre64):

250/7 is the per month interest ... so 12 of those equals a year

OpenStudy (anonymous):

35.71?

OpenStudy (amistre64):

that was the interest per month; and for 12 months how much would that have been?

OpenStudy (anonymous):

ok so then it would be $428.52

OpenStudy (amistre64):

lets avoid truncating; it builds errors

OpenStudy (amistre64):

12*(250/7) = 428.5714; is better soo: 3600, times interest = 428.5714 divide each side by 3600 to determine the interest

OpenStudy (amistre64):

well, both ways are equal to 4 decimal places so i spose thats fine as well

OpenStudy (anonymous):

I got 119047

OpenStudy (amistre64):

you missed your decimal .... but yeah

OpenStudy (amistre64):

so lets go with 11.9%

OpenStudy (anonymous):

I thought you said 4 decimal places?

OpenStudy (amistre64):

428.52/3600 = 0.1190333333 428.5714/3600 = 0.119047619 both values are equal up to 4 decimal places: 0.1190

OpenStudy (anonymous):

oh I see that makes logical sense. :) Thank you so much! :D I like when people show all the work so I can understand it verese giving me answers! :D

OpenStudy (amistre64):

youre welcome